### v1

@@ C.1 Approximating $C(K),\Sigma_{K}$ with finite horizon
We aim to show that, for $K\in\mathcal{S}$, it is possible to approximate both $C(K)$ and $\Sigma_{K}$ with any desired accuracy: {Lemma}[Approximating $C(K)$ and $\Sigma_{K}$ with finite horizon] For any $K\in\mathcal{S}$, let $\Sigma_{K}^{(l)}:=\frac{1}{l}\sum_{t=0}^{l-1}\mathbb{E}_{x_{0},w_{t}}\left[x_{t}x_{t}^{\top}\right]$ and $C^{(l)}(K):=\frac{1}{l}\sum_{t=0}^{l-1}\mathbb{E}_{x_{0},w_{t}}\left[x_{t}^{\top}Q_{K}x_{t}\right]$. If
@@
(64) $l\geq\frac{2C(K)}{\epsilon_{C}\lambda_{1}(\Sigma_{w})}\left(\frac{C(K)\lVert\Sigma_{0}\rVert}{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}+\frac{C^{2}(K)}{\lambda_{1}(Q)\lambda_{1}^{2}(\Sigma_{w})}+\frac{C(K)}{\lambda_{1}(Q)}\right),$
@@
and the operator $\mathcal{C}_{Y}(X):=\sum_{t=0}^{+\infty}Y^{t}XY^{t~{}\top}$ with $X\succ 0$ and $\lVert Y\rVert\leq 1$ From the definition of $\Sigma_{K}^{(l)}$:
$\begin{split}\Sigma_{K}^{(l)}&=\frac{1}{l}\left[\Sigma_{0}+\sum_{t=1}^{l-1}\left[(A_{K})^{t}\Sigma_{0}(A_{K})^{t~{}\top}+\sum_{k=0}^{t-1}(A_{K})^{t-k}\Sigma_{w}(A_{K})^{t-k~{}\top}\right]\right]\\ &=\frac{1}{l}\sum_{t=0}^{l-1}\left[(A_{K})^{t}\Sigma_{0}(A_{K})^{t~{}\top}\right]+\frac{1}{l}\sum_{t=1}^{l-1}\left[\sum_{k=0}^{t-1}(A_{K})^{t-k}\Sigma_{w}(A_{K})^{t-k~{}\top}\right]\\ &=\Sigma_{K}+\frac{(\mathcal{C}_{A_{K}}(\Sigma_{0}-P_{K}))-A_{K}^{l}(\mathcal{C}_{A_{K}}(\Sigma_{0})-\tilde{P}_{K}^{w})A_{K}^{l~{}\top}-2\Sigma_{K}}{l}\\ \end{split}$
@@
$\begin{split}\lVert\Sigma_{K}^{(l)}-\Sigma_{K}\rVert&\leq\frac{2\lVert\mathcal{C}_{A_{K}}(\Sigma_{0}-P_{K})\rVert+2\lVert\Sigma_{K}\rVert}{l}\\ &\leq\frac{2}{l}{\left(\frac{C(K)\lVert\Sigma_{0}\rVert}{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}+\frac{C^{2}(K)}{\lambda_{1}(Q)\lambda_{1}^{2}(\Sigma_{w})}+\frac{C(K)}{\lambda_{1}(Q)}\right)}=:\epsilon_{\Sigma}\end{split}$
@@
$\begin{split}C^{(l)}(K):&=\frac{1}{l}\sum_{t=0}^{l-1}\mathbb{E}\left[x_{t}^{\top}Q_{K}x_{t}\right]=\frac{1}{l}\sum_{t=0}^{l-1}\mathrm{Tr}\left[Q_{K}\Sigma_{t}\right]=\mathrm{Tr}\left[Q_{K}\frac{1}{l}\sum_{t=0}^{l-1}\Sigma_{t}\right]\\ &=\mathrm{Tr}\left[Q_{K}(\Sigma_{K}+\frac{(\mathcal{C}_{A_{K}}(\Sigma_{0})-\tilde{P}_{K}^{w})-A_{K}^{l}(\mathcal{C}_{A_{K}}(\Sigma_{0})-\tilde{P}_{K}^{w})A_{K}^{l~{}\top}-2\Sigma_{K}}{l})\right]\\ &=C(K)+\frac{1}{l}\mathrm{Tr}\left[{Q_{K}[(\mathcal{C}_{A_{K}}(\Sigma_{0})-\tilde{P}_{K}^{w})-A_{K}^{l}(\mathcal{C}_{A_{K}}(\Sigma_{0})-\tilde{P}_{K}^{w})A_{K}^{l~{}\top}-2\Sigma_{K}]}\right]\\ \end{split}$
@@
$\begin{split}|C^{(l)}(K)-C(K)|&\leq\frac{2C(K)}{l\lambda_{1}(\Sigma_{w})}\left(\frac{C(K)\lVert\Sigma_{0}\rVert}{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}+\frac{C^{2}(K)}{\lambda_{1}(Q)\lambda_{1}^{2}(\Sigma_{w})}+\frac{C(K)}{\lambda_{1}(Q)}\right)=:\epsilon_{C}\\ \end{split}$
@@
(65) $l\geq\frac{2C(K)}{\epsilon_{C}\lambda_{1}(\Sigma_{w})}\left(\frac{C(K)\lVert\Sigma_{0}\rVert}{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}+\frac{C^{2}(K)}{\lambda_{1}(Q)\lambda_{1}^{2}(\Sigma_{w})}+\frac{C(K)}{\lambda_{1}(Q)}\right).$
@@ C.2 Smoothing and the gradient descent analysis
(66) $C_{r}(K):=\mathbb{E}_{U\sim\mathbb{B}_{r}}\left[C(K+U)U\right],$
where $\mathbb{B}_{r}$ is the uniform distribution over all matrices with Frobenius norm less than $r$ (the entire ball) {Lemma}[Gradient of smoothing function]
(67) $\nabla C_{r}(K)=\frac{n_{x}n_{u}}{r^{2}}\mathbb{E}_{U\sim\mathbb{S}_{r}}\left[C(K+U)U\right],$
where $\mathbb{S}_{r}$ is the uniform distribution over all matrices with Frobenius norm $r$ (the sphere) The proof of Lemma 66 is given in [10, Lemma 26].
@@ C.3 Estimating $\nabla C(K)$ with finitely many infinite-horizon Rollouts
(68) ¯$\bar{\nabla}C(K):=\frac{1}{n}\sum^{n}_{k=1}\frac{n_{x}n_{u}}{r^{2}}C(K+U_{k})U_{k},$
and $U_{i}$ is uniformly distributed and $\lVert U_{i}\rVert_{F}=r$. {Lemma} Given an accuracy $\epsilon$ and probability $\delta$, which can be decomposed by $\epsilon_{1}$ and $\epsilon_{2}$ ($\epsilon=\epsilon_{1}+\epsilon_{2}$), there exist an exploration radius $r_{\max}$ which is a function of $\epsilon_{1}$ and the number of rollouts $n$ which is a function of $\epsilon_{1},\epsilon_{2}$ such that when $r\leq r_{\max}$ and $n\geq N$, with high probability (at least $1-\delta$),
(69) ¯$\lVert\bar{\nabla}C(K)-\nabla C(K)\rVert\leq\epsilon$
{Proof}
First, we bridge ¯$\bar{\nabla}C(K)$ and $\nabla C(K)$ with $\nabla C_{r}(K)$,
¯$\bar{\nabla}C(K)-\nabla C(K)=(\nabla C_{r}(K)-\nabla C(K))+(\bar{\nabla}C(K)-\nabla C_{r}(K))$
For the first term, let $r_{\max}\leq\min\{h(C(K)),\lVert K\rVert\}$,we have $C(K+U_{k})-C(K)\leq h_{C}(C(K))\lVert U_{k}\rVert_{F}$, Then we can get $|C(K+U_{k})|\leq C(K)+h_{C}(C(K))\lVert U_{k}\rVert_{F}$. From Lemma 2.2, if
(70) $r_{\max}\leq\min\{h(C(K)),\lVert K\rVert,\frac{\epsilon_{1}}{h_{\nabla}(C(K))}\},$
@@
$\lVert\nabla C(K+U_{k})-{\nabla}C(K)\rVert\leq{\epsilon_{1}}.$
@@
(71) $\lVert\nabla C_{r}(K)-\nabla C(K)\rVert\leq{\epsilon_{1}}.$
For the second term ¯$\lVert\bar{\nabla}C(K)-\nabla C_{r}(K)\rVert$, we define the sample $Z_{k}:=\frac{n_{x}n_{u}}{r^{2}}C(K+U_{k})U_{k}$
@@
$\lVert Z\rVert=\lVert\nabla C_{r}(K)\rVert\leq\lVert\nabla C(K)\rVert+{\epsilon_{1}}\leq{\epsilon_{1}}+b_{\nabla}(C(K)).$
@@
(72) $\lVert Z_{k}-Z\rVert\leq\lVert Z_{k}\rVert+\lVert Z\rVert\leq\underbrace{\frac{n_{x}n_{u}\left(C(K)+rh_{C}(C(K))\right)}{r}+\epsilon_{1}+b_{\nabla}(C(K))}_{=:\alpha_{4}(C(K),\epsilon_{1})}$
@@
(73) $\begin{split}&~{}~{}~{}~{}\lVert\mathbb{E}(Z_{k}Z_{k}^{\top})-ZZ^{\top}\rVert=\lVert\mathbb{E}(Z_{k}^{\top}Z_{k})-Z^{\top}Z\rVert\\ &\leq\lVert\mathbb{E}(Z_{k}^{\top}Z_{k})\rVert_{F}+\lVert Z^{\top}Z\rVert_{F}\leq\max_{Z_{k}}(\lVert Z_{k}\rVert_{F})^{2}+\lVert Z\rVert^{2}_{F}\\ &\leq\underbrace{\max\{n_{x},n_{u}\}^{2}\left(\frac{n_{x}n_{u}\left(C(K)+rh_{C}(C(K))\right)}{r}\right)^{2}+\left(\epsilon_{1}+b_{\nabla}(C(K))\right)^{2}}_{=:\alpha_{5}(C(K),\epsilon_{1})}\end{split}$
We aim to bound ¯$\lVert\bar{\nabla}C(K)-\nabla C_{r}(K)\rVert\leq\epsilon_{2}$, so the number of rollouts much be greater and equal than $N$ (Matrix Bernstein Inequality, [11, Lemma B.5])
(74) $N:=\frac{2\min\{n_{x},n_{u}\}}{\epsilon_{2}^{2}}\left(\alpha_{4}^{2}(C(K),\epsilon_{1})+\frac{\alpha_{5}(C(K),\epsilon_{1})\epsilon_{2}}{3\sqrt{\min(n_{x},n_{u})}}\right)\log\left[\frac{n_{x}+n_{u}}{\delta}\right],$
@@
[¯$\mathbb{P}\left[\lVert\bar{\nabla}C(K)-\nabla C_{r}(K)\rVert_{F}\leq\epsilon_{2}\right]\geq 1-\delta.$
We sum up all the terms and get the following if $r\geq r_{\max}$ and ${n}\geq N$, where $r_{min}$ and $N$ are defined in (70) and (74), we have
(75) [¯$\mathbb{P}\left[\lVert\bar{\nabla}C(K)-\nabla C(K)\rVert_{F}\leq\epsilon\right]\geq 1-\delta.$
@@
The noise $w_{t}$ is a Gaussian distributed vector with zero mean and covariance $\Sigma_{w}$. Let $\Sigma_{w}=V\Lambda V^{\top}$ be the eigenvalue decomposition, where $V$ is a diagonal matrix with diagonal entries $\sigma_{1},...,\sigma_{n_{x}}$. Then we have the vector $W_{t}=V^{-1}w_{t}$, with zero mean and variance $\Lambda$. By the Chebyshev’s inequality [32, Corollary 1.2.5], we have
(76) $\mathbb{P}\left\{\lVert w_{t}\rVert\leq\bar{w}\right\}=\mathbb{P}\left\{\lVert W_{t}\rVert\leq\bar{w}\right\}\geq\left\{1-\frac{n_{x}\lambda_{1}(\Sigma_{w})}{\bar{w}^{2}}\right\}^{n_{x}}=:\delta_{w}({\bar{w}})$
Based on (76), we can derive the following Lemma: {Lemma}[Bound of state $\lVert x_{t}\rVert$] For any initial state $\lVert x_{0}\rVert\leq L_{0}$ and $K\in\mathcal{S}$, given a desired probability $\delta$, the state response $\lVert x_{t}\rVert$ expressed as (3) is bounded by $\bar{L}:=L_{0}+t\left(\frac{\lambda_{1}(\Sigma_{w})^{n_{x}}}{1-\delta^{\frac{1}{t}}}\right)^{\frac{1}{2n_{x}}}$ with probability at least $\delta$. {Proof} From (3), we can get:
@@
If for all $t\in[0,t-1]$,$\lVert w_{t}\rVert\leq\bar{w}$, then $\lVert x_{t}\rVert\leq L_{0}+t\bar{w}=\bar{L}$. Then the probability of $\lVert x_{t}\rVert\leq\bar{L}(\bar{w})$ is at least $\delta:=\delta_{w}(\bar{w})^{t}$. Together with (76), we can then derive $\bar{w}=\sqrt{\frac{n_{x}\lambda_{1}(\Sigma_{w})}{1-\delta^{\frac{1}{tn_{x}}}}}$. We conclude the proof.

### v2

@@ C.1 Approximating $C(K),\Sigma_{K}$ with finite horizon
We aim to show that, for $K\in\mathcal{S}$, it is possible to approximate both $C(K)$ and $\Sigma_{K}$ with any desired accuracy:
Lemma 6 (Approximating $C(K)$ and $\Sigma_{K}$ with finite horizon)
For any $K\in\mathcal{S}$, let
$\Sigma_{K}^{(l)}:=\frac{1}{l}\sum_{t=0}^{l-1}\mathbb{E}_{x_{0},w_{t}}\left[x_{t}x_{t}^{\top}\right]~\mathrm{and}~C^{(l)}(K):=\frac{1}{l}\sum_{t=0}^{l-1}\mathbb{E}_{x_{0},w_{t}}\left[x_{t}^{\top}Q_{K}x_{t}\right].$ (64)
Given any desired accuracy $\epsilon_{\Sigma},\epsilon_{C}>0$, if
@@
$l\geq\frac{2C(K)}{\epsilon_{C}\lambda_{1}(\Sigma_{w})}\left(\frac{C(K)\lVert\Sigma_{0}\rVert}{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}+\frac{C^{2}(K)}{\lambda_{1}(Q)\lambda_{1}^{2}(\Sigma_{w})}+\frac{C(K)}{\lambda_{1}(Q)}\right),$ (65)
@@
and the operator $\mathcal{C}_{Y}(X):=\sum_{t=0}^{+\infty}Y^{t}XY^{t~\top}$ with $X\succ 0$ and $\lVert Y\rVert\leq 1$ . From the definition of $\Sigma_{K}^{(l)}$:
[…]
@@
$\begin{split}\lVert\Sigma_{K}^{(l)}-\Sigma_{K}\rVert&\leq\frac{2\lVert\mathcal{C}_{A_{K}}(\Sigma_{0}-P_{K})\rVert+2\lVert\Sigma_{K}\rVert}{l}\\ &\leq\frac{2}{l}{\left(\frac{C(K)\lVert\Sigma_{0}\rVert}{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}+\frac{C^{2}(K)}{\lambda_{1}(Q)\lambda_{1}^{2}(\Sigma_{w})}+\frac{C(K)}{\lambda_{1}(Q)}\right)}=:\epsilon_{\Sigma}.\end{split}$
@@
[…]
@@
[…]
@@
$l\geq\frac{2C(K)}{\epsilon_{C}\lambda_{1}(\Sigma_{w})}\left(\frac{C(K)\lVert\Sigma_{0}\rVert}{\lambda_{1}(Q)\lambda_{1}(\Sigma_{w})}+\frac{C^{2}(K)}{\lambda_{1}(Q)\lambda_{1}^{2}(\Sigma_{w})}+\frac{C(K)}{\lambda_{1}(Q)}\right).$ (66)
@@ C.2 Smoothing and the gradient descent analysis
$C_{r}(K):=\mathbb{E}_{U\sim\mathbb{B}_{r}}\left[C(K+U)U\right],$ (67)
where $\mathbb{B}_{r}$ is the uniform distribution over all matrices with Frobenius norm less than $r$ (the entire ball)
Lemma 7 (Gradient of smoothing function)
$\nabla C_{r}(K)=\frac{n_{x}n_{u}}{r^{2}}\mathbb{E}_{U\sim\mathbb{S}_{r}}\left[C(K+U)U\right],$ (68)
where $\mathbb{S}_{r}$ is the uniform distribution over all matrices with Frobenius norm $r$ (the sphere)
The proof of Lemma 7 is given in [9, Supplementary Material, Lemma 26].
@@ C.3 Estimating $\nabla C(K)$ with finitely many infinite-horizon Rollouts
$\bar{\nabla}C(K):=\frac{1}{n}\sum^{n}_{k=1}\frac{n_{x}n_{u}}{r^{2}}C(K+U_{k})U_{k},$ (69)
and $U_{i}$ is uniformly distributed and $\lVert U_{i}\rVert_{F}=r$.
Lemma 8
Given an accuracy $\epsilon_{nr}$ and probability $\delta_{n}$, which can be expressed by $\epsilon_{r}$ and $\epsilon_{n}$ ($\epsilon_{nr}=\epsilon_{r}+\epsilon_{n}$), there exist an radius $r_{\max}$ which is a function of $\epsilon_{r}$ and the number of rollouts $N_{1}$ which is a function of $\epsilon_{r},\epsilon_{n}$ such that when $r\leq r_{\max}$ and $n\geq N_{1}$,
$\mathbb{P}\left\{\lVert\bar{\nabla}C(K)-\nabla C(K)\rVert\leq\epsilon_{nr}\right\}\geq 1-\delta_{n}.$ (70)
Proof 9
First, we bridge $\bar{\nabla}C(K)$ and $\nabla C(K)$ with $\nabla C_{r}(K)$,
$\bar{\nabla}C(K)-\nabla C(K)=(\nabla C_{r}(K)-\nabla C(K))+(\bar{\nabla}C(K)-\nabla C_{r}(K)).$
For the first term, let $r_{\max}\leq\min\{h(C(K)),\lVert K\rVert\}$,we have $C(K+U_{k})-C(K)\leq h_{C}(C(K))\lVert U_{k}\rVert_{F}$, Then we can get $|C(K+U_{k})|\leq C(K)+h_{C}(C(K))\lVert U_{k}\rVert_{F}$. From Lemma 5, if
$r_{\max}\leq\min\{h(C(K)),\lVert K\rVert,\frac{\epsilon_{r}}{h_{\nabla}(C(K))}\},$ (71)
@@
$\lVert\nabla C(K+U_{k})-{\nabla}C(K)\rVert\leq{\epsilon_{r}}.$
@@
$\lVert\nabla C_{r}(K)-\nabla C(K)\rVert\leq{\epsilon_{r}}.$ (72)
For the second term $\lVert\bar{\nabla}C(K)-\nabla C_{r}(K)\rVert$, we define the sample $Z_{k}:=\frac{n_{x}n_{u}}{r^{2}}C(K+U_{k})U_{k}$
@@
$\lVert Z\rVert=\lVert\nabla C_{r}(K)\rVert\leq\lVert\nabla C(K)\rVert+{\epsilon_{r}}\leq{\epsilon_{r}}+b_{\nabla}(C(K)).$
@@
$\lVert Z_{k}-Z\rVert\leq\lVert Z_{k}\rVert+\lVert Z\rVert\leq\underbrace{\frac{n_{x}n_{u}\left(C(K)+rh_{C}(C(K))\right)}{r}+\epsilon_{r}+b_{\nabla}(C(K))}_{=:\alpha_{4}(C(K),\epsilon_{r})}$ (73)
@@
$\begin{split}&~~~~\lVert\mathbb{E}(Z_{k}Z_{k}^{\top})-ZZ^{\top}\rVert=\lVert\mathbb{E}(Z_{k}^{\top}Z_{k})-Z^{\top}Z\rVert\\ &\leq\lVert\mathbb{E}(Z_{k}^{\top}Z_{k})\rVert_{F}+\lVert Z^{\top}Z\rVert_{F}\leq\max_{Z_{k}}(\lVert Z_{k}\rVert_{F})^{2}+\lVert Z\rVert^{2}_{F}\\ &\leq\underbrace{\max\{n_{x},n_{u}\}^{2}\left(\frac{n_{x}n_{u}\left(C(K)+rh_{C}(C(K))\right)}{r}\right)^{2}+\left(\epsilon_{r}+b_{\nabla}(C(K))\right)^{2}}_{=:\alpha_{5}(C(K),\epsilon_{r})}\end{split}$ (74)
We aim to bound $\lVert\bar{\nabla}C(K)-\nabla C_{r}(K)\rVert\leq\epsilon_{n}$, so the number of rollouts must be greater or equal to $N_{1}$ (from Matrix Bernstein Inequality, [10, Lemma B.5])
$N_{1}:=\frac{2\min\{n_{x},n_{u}\}}{\epsilon_{n}^{2}}\left(\alpha_{4}^{2}(C(K),\epsilon_{r})+\frac{\alpha_{5}(C(K),\epsilon_{r})\epsilon_{n}}{3\sqrt{\min(n_{x},n_{u})}}\right)\log\left[\frac{n_{x}+n_{u}}{\delta}\right],$ (75)
@@
$\mathbb{P}\left[\lVert\bar{\nabla}C(K)-\nabla C_{r}(K)\rVert_{F}\leq\epsilon_{n}\right]\geq 1-\delta_{n}.$
We sum up all the terms and get the following: if $r\geq r_{\max}$ and ${n}\geq N$, where $r_{min}$ and $N_{1}$ are defined in (71) and (75), we have
$\mathbb{P}\left[\lVert\bar{\nabla}C(K)-\nabla C(K)\rVert_{F}\leq\epsilon_{nr}\right]\geq 1-\delta_{n}.$ (76)
@@
The noise $w_{t}$ is a Gaussian-distributed vector with zero mean and covariance $\Sigma_{w}$. Let $\Sigma_{w}=V\Lambda V^{\top}$ be the eigenvalue decomposition, where $\Lambda$ is a diagonal matrix with diagonal entries $\sigma_{1},...,\sigma_{n_{x}}$. Then we have the vector $W_{t}=V^{-1}w_{t}$, with zero mean and variance $\Lambda$. By the Chebyshev’s inequality [52, Corollary 1.2.5], we have
$\mathbb{P}\left\{\lVert w_{t}\rVert\geq\bar{w}\right\}=\mathbb{P}\left\{\lVert W_{t}\rVert\geq\bar{w}\right\}\leq\frac{\sum_{k=1}^{n_{x}}\sigma_{k}}{\bar{w}^{2}}=\frac{\mathrm{Tr}(\Sigma_{w})}{\bar{w}^{2}}=:\delta_{w}({\bar{w}})$ (77)
Based on (77), we can derive the following Lemma:
Lemma 9 (Bound of state $\lVert x_{t}\rVert$)
For any initial state $\lVert x_{0}\rVert\leq L_{0}$ and $K\in\mathcal{S}$, given a desired probability $\delta_{x}$, the state response $\lVert x_{t}\rVert$ expressed as (3) is bounded by $\bar{L}:=L_{0}+t\frac{\mathrm{Tr(\Sigma_{w})}}{1-(1-\delta_{x})^{1/t}}$ with probability at least $1-\delta_{x}$.
Proof 10
From (3), we can get:
@@
If for all $t\in[0,t-1]$$\lVert w_{t}\rVert\leq\bar{w}$, then $\lVert x_{t}\rVert\leq L_{0}+t\bar{w}=\bar{L}$. Then the probability of $\lVert x_{t}\rVert\leq\bar{L}$ is at least $1-\delta=(1-\delta_{w}(\bar{w}))^{t}$. Together with (77), we can then derive $\bar{w}=\frac{\mathrm{Tr(\Sigma_{w})}}{1-(1-\delta_{x})^{1/t}}$. We conclude the proof.